OpenStudy (anonymous):

PLEASE HELP Solve x^2-8x=20

OpenStudy (anonymous):

half of 8 is 4, and 4 squared is 16 so start with \[(x-4)^2=20+16\] and go form there

OpenStudy (anonymous):

\[(x-4)^2=36\] \[x-4=\pm\sqrt{36}\] \[x-4=6\] or \[x-4=-6\] one more step and you are done

OpenStudy (anonymous):

x=10 and x=-2

OpenStudy (anonymous):

I got that thanksss

OpenStudy (anonymous):

No problem